### v2

# Theoretical study on the crystal structure of a bilayer nickel-oxychloride Sr3Ni2O5Cl2 and analysis on the occurrence of possible unconventional superconductivity

Abstract
The discovery of superconductivity under high pressure with $T_{c}$ exceeding 80 K in a bilayer nickelate La3Ni2O7 has led to a strong desire to realize similar high $T_{c}$ phenomena at ambient pressure. As one possible path toward realizing superconductivity at ambient pressure, we here propose to consider Sr3Ni2O5Cl2 as a possible candidate. In this study, we theoretically investigate the electronic structure of Sr3Ni2O5Cl2 and its structural stability. Our phonon calculation shows that this compound with the $I4/mmm$ tetragonal structure is dynamically stable even at ambient pressure. The characteristic crystal field in this compound lowers the Ni-$d_{3z^{2}-r^{2}}$ orbital energy, by which the Ni-$d_{3z^{2}-r^{2}}$ orbital becomes rather closer to the half-filling in Sr3Ni2O5Cl2 than La3Ni2O7. As a result, we find that superconductivity is enhanced even though a relatively strong orbital hybridization between the $t_{2g}$ and $e_{g}$ orbitals is somewhat detrimental for superconductivity. We also check the formation enthalpy, which shows that the high-pressure synthesis can be a good way to actually produce Sr3Ni2O5Cl2. We find that Sr3Ni2O5Cl2 is a promising new candidate of bilayer-nickelate superconductors, which can possess even higher $T_{c}$ than pressurized La3Ni2O7, at ambient pressure.
I Introduction
Discovery of superconductivity with a maximum $T_{c}$ of above 80 K in a bilayer Ruddlesden-Popper nickelate La3Ni2O7 (Fig. 1(a)) under pressure has sparked enormous interest in the field of condensed matter physics [1]. Although under pressure, $T_{c}$ exceeding the liquid nitrogen boiling temperature, and also the origin of superconductivity likely being unconventional [2, 3, 4], have led to a huge amount of experimental and theoretical studies after the discovery [1, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 3, 4, 69, 2, 70, 71, 72, 73, 74, 75, 76, 77]. In fact, the possibility of superconductivity in this material was discussed theoretically even before its experimental discovery by some of the previous authors [78], but one viewpoint that was missing at that time was the precise crystal structure, namely, although a tetragonal symmetry was simply assumed in Ref. 78, the actual structural symmetry is orthorhombic at ambient pressure. There are now accumulating studies suggesting that tetragonal symmetry of the crystal structure may be important for the occurrence of superconductivity in La3Ni2O7 [32, 42, 41], which is accomplished by applying pressure. One speculation regarding the reason for this can be that orthorhombicity favors density waves, which suppresses superconductivity. Interestingly, also for the trilayer compound La4Ni3O10, the superconducting phase appears by applying a sufficient pressure and the tetragonal structure is realized [31, 38, 39, 40, 48]. Then, a natural motivation arises for searching materials that have tetragonal symmetry at ambient pressure, and also have electronic properties similar to that of La3Ni2O7 at high pressures.
[…]
In this study, we theoretically investigate the electronic structure of Sr3Ni2O5Cl2 and its structural stability. Our phonon calculation shows that this compound with the $I4/mmm$ tetragonal structure is dynamically stable, i.e., does not have imaginary-frequency phonon modes, even at ambient pressure. The Ni-$d_{3z^{2}-r^{2}}$ orbital energy is pushed down by the crystal field where half of the apical O2- ions are replaced with a Cl- ion. By this effect, the Ni-$d_{3z^{2}-r^{2}}$ orbital becomes rather closer to the half-filling in Sr3Ni2O5Cl2 than La3Ni2O7 whereas the stoichiometric Ni-$3d$ band filling are decreased in Sr3Ni2O5Cl2 ($d^{7}$) compared with La3Ni2O7 ($d^{7.5}$). We also investigate superconductivity in this compound with the multi-orbital Hubbard model constructed from our first-principles calculation, using the fluctuation exchange (FLEX) approximation combined with the linearized Eliashberg equation. While orbital hybridization between the $t_{2g}$ and $e_{g}$ orbitals, which is enhanced in this characteristic crystal field, is detrimental for superconductivity, a nearly half-filled Ni-$d_{3z^{2}-r^{2}}$ orbital enhances superconductivity. In total, we find that Sr3Ni2O5Cl2 is a promising new candidate of bilayer-nickelate superconductors, which can possess even higher $T_{c}$ than pressurized La3Ni2O7, at ambient pressure. We also check the formation enthalpy for synthesis of this compound. Our calculation shows that the high-pressure synthesis can be a good way to experimentally produce Sr3Ni2O5Cl2.
Figure 1: (a) Crystal structure of La3Ni2O7 at high pressure. (b) Crystal structure and (c) phonon band dispersion for Sr3Ni2O5Cl2 without applying pressure. Panels (b) and (c) are shown for the space group of $I4/mmm$. Crystal structures in this paper were depicted using the VESTA software [82]. Phonon dispersions in this paper were depicted in the Brillouin zone of the conventional tetragonal lattice.
II Methods
For first-principles calculation, we used the PBEsol exchange-correlation functional [83] and the projector augmented wave method as implemented in Vienna ab initio Simulation Package (VASP) [84, 85, 86, 87, 88]. Core-electron states in PAW potentials were [Kr]$4d^{10}$, [Ar]$3d^{10}$, [Ar], [Ne], and [He] for La, Sr, Ni, Cl, and O, respectively. We used a plane-wave cutoff energy of 600 eV for Kohn-Sham orbitals without including the spin-orbit coupling for simplicity. A $12\times 12\times 12$ ${\bm{k}}$-mesh was used for self-consistent-field calculations.
We performed structural optimization until the Hellmann-Feynman force becomes less than 0.01 eV Å-1 for each atom. To check the stability of the tetragonal ($I4/mmm$) phase, we calculated the phonon dispersion for several materials. For this purpose, we used the finite displacement method as implemented in the Phonopy [89, 90] software in combination with VASP. We used a $4\times 4\times 1$ ${\bm{q}}$-mesh for a conventional tetragonal unit cell. For a $4\times 4\times 1$ supercell used for finite-displacement calculations, we used a $3\times 3\times 2$ ${\bm{k}}$-mesh.
[…]
We adopted FLEX [94, 95] to analyze electron correlation effects for the Hubbard model as was done in Ref. [8]. We calculated the self-energy induced by the spin-fluctuation formulated as shown in the literatures [96, 97, 98] in a self-consistent calculation. The real part of the self-energy at the lowest Matsubara frequency was subtracted in the same manner as Ref. [99] to maintain the band structure around the Fermi level obtained by first-principles calculation. We used the linearized Eliashberg equation to study the possibility of superconductivity in Sr3Ni2O5Cl2 comparing with La3Ni2O7. The renormalized Green’s functions obtained by FLEX were plugged into this equation. Note that the pairing interaction kernel in this equation was obtained from the FLEX Green’s function as a purely electronic one (i.e. phonon-mediated pairing interaction was not considered), which is mainly dominated by spin fluctuations in the present case. Since the eigenvalue $\lambda$ of the Eliashberg equation monotonically increases upon lowering the temperature $T$, and reaches unity at $T=T_{c}$, we adopted $\lambda$ calculated at a fixed temperature, $T=0.01$ eV, as a measure of superconductivity. For convenience, we will call the eigenfunction (with the largest eigenvalue) of the linearized Eliashberg equation at the lowest Matsubara frequency $i\omega$ $=$ $i\pi k_{\rm B}T$ the “superconducting gap function”. We took a 16$\times$16$\times$4 ${\bm{k}}$-mesh and 2048 Matsubara frequencies for the FLEX calculation.
For calculating the formation energy, we adopted space groups taken from experimental studies: $Fm{\bar{3}}m$ for SrO, SrCl2, Ni, NiO, $I4/mmm$ for SrCl2, Sr2NiO2Cl2 [100], $Cmcm$ for SrNiO2 [101], and $P6_{3}/mmc$ for SrNiO3 [102]. An isolated O2 molecule was calculated in a 15 Å $\times$ 15 Å $\times$ 15 Å cell considering the spin-triplet ground state with the spin-polarized calculation.
III Results and Discussions
III.1 Electronic and phonon band structures for Sr3Ni2O5Cl2
As described in the Introduction, a key idea in this study is to employ a mixed-anion strategy [103] for stabilizing the tetragonal $I4/mmm$ phase without applying the pressure. In fact, for Sr3Ni2O5Cl2, while the NiO2 plane is somewhat buckled; the equatorial Ni-O-Ni angle is 162 as shown in Fig. 1(b), we found that the tetragonal structure ($a=3.779$, $c=24.532$) is dynamically stable as shown by our phonon calculation (Fig. 1(c)) where no imaginary-frequency modes appear without applying the pressure. This is in sharp contrast to the case of La3Ni2O7, where the tetragonal structure is unstable at ambient pressure and high pressure is required to stabilize it [32, 42, 41]. Our phonon calculations for La3Ni2O7 also support this feature as shown in Fig. 6, where the imaginary-frequency modes appear for 0 GPa while they are absent for 20 GPa. Although we also checked the stability of the tetragonal structure for Sr3Ni2O4Cl3, where all the apical oxygens were replaced with chlorines, we found that the tetragonal structure was unstable for this case as shown in Fig. 7. We consider that one important difference among these compounds is the average valence number of the nickel atom: Ni2.5+ ($d^{7.5}$) for La3Ni2O7 and Sr3Ni2O4Cl3 while Ni3+ ($d^{7}$) for Sr3Ni2O5Cl2. Here, the size of the nickel ion becomes smaller in Sr3Ni2O5Cl2, which in general suppresses the rotational and tilting instability of octahedra in Ruddlesden-Popper phase materials.
It is noteworthy that the mixed-anion strategy drastically changes the crystal field of a nickel ion. Figure 2 presents the electronic band structures and partial density of states (pDOS) for La3Ni2O7 under pressure of 20 GPa and Sr3Ni2O5Cl2 without applying the pressure shown with an orbital weight of each Ni-$d$ orbital. We summarize important differences between two materials from the viewpoint of the crystal field as follows.
As mentioned above, the average valence number of the nickel atom is different: Ni2.5+ ($d^{7.5}$) for La3Ni2O7 and Ni3+ ($d^{7}$) for Sr3Ni2O5Cl2. Thus, the number of $d$ electrons for Sr3Ni2O5Cl2 is less than that for La3Ni2O7. Nevertheless, the Ni-$d_{3z^{2}-r^{2}}$ orbital is rather closer to the half-filling in Sr3Ni2O5Cl2 as shown in Figs. 2(b)(h). This is because of the characteristic crystal field for the nickel atom in Sr3Ni2O5Cl2, where a half of apical O2- ions is replaced with a Cl- ion, which pushes down the orbital energy of the Ni-$d_{3z^{2}-r^{2}}$ orbital. As a result, the number of $d_{x^{2}-y^{2}}$ electrons is decreased in Sr3Ni2O5Cl2. On the other hand, Ni-$d_{xy/yz/zx}$ orbitals, which we hereafter call them the $t_{2g}$ orbitals while it is not a rigorous term for this symmetry, raise their energy levels relative to the Ni-$d_{3z^{2}-r^{2}}$ orbital. This is because the present crystal field strongly pushes down the Ni-$d_{3z^{2}-r^{2}}$ orbital level as we have stated above. Consequently, the strong DOS peak of the $t_{2g}$ orbitals lies just below the Fermi energy. In addition, the crystal field is strongly asymmetric with respect to the $xy$ plane (i.e., $+z$ and $-z$ are strongly inequivalent), which allows a sizable hybridization among $e_{g}$ ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) and $t_{2g}$ orbitals. In the next section, we will discuss how these characteristics can affect superconductivity in this system.
Figure 2: Electronic band dispersion with a colored weight of Ni- (a) $d_{x^{2}-y^{2}}$, (b) $d_{3z^{2}-r^{2}}$, (c) $d_{xy}$, (d) $d_{yz}$, and (e) $d_{zx}$ orbitals, respectively, and (f) pDOS for La3Ni2O7 under pressure of 20 GPa. (g)–(l) Same plots for Sr3Ni2O5Cl2 without applying the pressure.
III.2 Model construction
In this and the next sections, we compared Sr3Ni2O5Cl2 without applying the pressure and La3Ni2O7 under pressure of 20 GPa, both of which have the $I4/mmm$ tetragonal structure in our calculation. We constructed a ten-orbital model consisting of five $3d$ orbitals centered at two Ni sites per unit cell. Figure 3(a) shows superposed band structures given by first-principles calculation and Wannier interpolation, where precise fitting around the Fermi level was achieved. Important parameter values extracted by Wannierization are given in Table 1, compared with those of four-orbital models consisting only of $e_{g}$ orbitals. The smaller differences in extracted parameters between the ten- and four-orbital models for La3Ni2O7 originate from the fact that the $e_{g}$ and $t_{2g}$ bands are well separated compared with Sr3Ni2O5Cl2, where the $t_{2g}$ and $e_{g}$ bands are rather entangled, as shown in Fig. 2. We found a relatively large hopping amplitude between $t_{2g}$ and $e_{g}$ orbitals: e.g., the nearest-neighbor intralayer hopping along the $x$ direction between the $d_{x^{2}-y^{2}}$ and $d_{xz}$ orbitals and that between the $d_{3z^{2}-r^{2}}$ and $d_{xz}$ orbitals are $-0.296$ and $0.199$ eV, respectively. This is due to the strongly asymmetric crystal field for a NiO5Cl octahedron as mentioned in the previous section. It is also noteworthy that the relatively strong buckling of NiO2 plane in Sr3Ni2O5Cl2 seems to enhance he vertical interlayer hopping $t_{\perp}$ between the $d_{3z^{2}-r^{2}}$ orbitals while the nearest-neighbor intralayer hoppings $t_{3z^{2}-r^{2}}$, $t_{x^{2}-y^{2}}$, and $t_{x^{2}-y^{2}{\text{-}}3z^{2}-r^{2}}$ are somewhat suppressed as shown in Table 1. For Sr3Ni2O5Cl2, we also found that the onsite energies of the $d_{xy}$ and $d_{yz/zx}$ orbitals relative to that for the $d_{3z^{2}-r^{2}}$ orbital are $-0.811$ eV and $-0.840$ eV, respectively, the size of which are comparable to $\Delta E=E_{x^{2}-y^{2}}-E_{3z^{2}-r^{2}}=0.776$ eV listed in Table 1.
Table 1: The orbital level offset $\Delta E=E_{x^{2}-y^{2}}-E_{3z^{2}-r^{2}}$ between the $d_{x^{2}-y^{2}}$ and the $d_{3z^{2}-r^{2}}$ orbitals, the vertical interlayer hopping $t_{\perp}$ between the $d_{3z^{2}-r^{2}}$ orbitals, and the nearest-neighbor intralayer hoppings $t_{3z^{2}-r^{2}}$, $t_{x^{2}-y^{2}}$, and $t_{x^{2}-y^{2}{\text{-}}3z^{2}-r^{2}}$ are displayed. The first column indicates the chemical formulae of the calculated materials, where the number in the round brackets indicate the number of orbitals.
(eV) $\Delta E$ $t_{\perp}$ $t_{3z^{2}-r^{2}}$ $t_{x^{2}-y^{2}}$ $t_{x^{2}-y^{2}{\text{-}}3z^{2}-r^{2}}$
Sr3Ni2O5Cl2 (10) $0.776$ $-0.744$ $0.041$ $-0.425$ $-0.086$
Sr3Ni2O5Cl2 (4) $0.920$ $-0.786$ $0.037$ $-0.383$ $-0.074$
La3Ni2O7 (10) $0.343$ $-0.660$ $-0.126$ $-0.499$ $-0.253$
La3Ni2O7 (4) $0.328$ $-0.659$ $-0.128$ $-0.501$ $-0.253$
In Fig. 3(b), we plot the number of electrons for each orbital per Ni against the hypothetically varied total band filling, assuming a rigid band. In this paper, the total band filling $n$ is defined as the number of electrons per spin in the primitive unit cell, e.g., $0\leq n\leq 10$ for ten-orbital models. The doping ratio $\Delta n$ is defined as the deviation of the total band filling from its value at stoichiometry. In Sr3Ni2O5Cl2, the number of electrons occupying the $d_{3z^{2}-r^{2}}$ band per Ni, called $n[d_{3z^{2}-r^{2}}]$ hereafter, exceeds unity (half-filled) already at stoichiometry, consistent with the observation given in Section III.1. Also, the $d_{yz/zx}$ orbitals are not fully filled due to the hybridization with the $e_{g}$ orbitals. Contrastingly, in La3Ni2O7, the $d_{yz/zx}$ orbitals are fully filled in the whole doping range suggesting their irrelevance for superconductivity and justifying the previously proposed four-orbital model [8].
Based on the observation described in this section, we used the ten-orbital model for Sr3Ni2O5Cl2 and the four-orbital model (i.e., $e_{g}$ model) for La3Ni2O7 at 20 GPa in the following analysis of superconductivity.
Figure 3: (a) Band structure of the ten-orbital model (pink solid lines) superposed on that obtained from first-principles calculation (black dotted lines) for Sr3Ni2O5Cl2 (left) and those for the ten-orbital model (middle) and the four-orbital model (right) of La3Ni2O7, (b) the number of electrons per Ni for each orbital as a function of doping ratio $\Delta n$ for Sr3Ni2O5Cl2 (left) and La3Ni2O7 (right), (c) the eigenvalue $\lambda$ of the linearized Eliashberg equation at $T=0.01$ eV as a function of $\Delta n$ (left) and $n[d_{3z^{2}-r^{2}}]$ (right), where the case of stoichiometry is indicated by arrows. In panel (c), the legend (i) indicates the case that the $t_{2g}-e_{g}$ interorbital interactions are turned off, and (ii) all the interorbital interactions are turned off. The legend “La3Ni2O7” in panel (c) indicates the result of four-orbital-model calculation.
III.3 FLEX analysis on superconductivity in Sr3Ni2O5Cl2
In the left panel of Fig. 3(c), we plot the eigenvalue $\lambda$ of the linearized Eliashberg equation against $\Delta n$. For comparison, we also plot $\lambda$ for models where (i) the $t_{2g}-e_{g}$ interorbital interactions are turned off, and (ii) all the interorbital interactions are turned off. $\lambda$ is also plotted for the four-orbital model of La3Ni2O7. By comparing $\lambda$ calculated with the original (full) model and that calculated by the condition (i), it can be seen that the $t_{2g}-e_{g}$ interorbital interactions somewhat suppress superconductivity in Sr3Ni2O5Cl2. Still, quite interestingly, the value of $\lambda$ of the full model of Sr3Ni2O5Cl2 is significantly larger than that of La3Ni2O7 at stoichiometry, suggesting the possibility that the material possesses a potential of even higher $T_{c}$ than 80 K at ambient pressure. Also, a slight electron doping may result in an even higher $T_{c}$. Electron doping is expected to be possible by partially substituting Sr by La, for which we have also verified the stability of the tetragonal symmetry as shown in Fig. 8(b), where no imaginary-frequency phonon modes appear for La0.3Sr2.7Ni2O5Cl2 in our first-principles calculation. On the other hand, hole doping is not good for superconductivity.
In Fig. 4, we plot the gap function in the band representation for Sr3Ni2O5Cl2 ($\Delta n=0.01$) and La3Ni2O7 ($\Delta n=0$). For Sr3Ni2O5Cl2, the gap function of the $d_{3z^{2}-r^{2}}$ portion of the band is large with a basically $s\pm$-wave structure, i.e., -- sign in the bonding and $+$ sign in the antibonding band, as in La3Ni2O7. The gap function of the $d_{x^{2}-y^{2}}$ portion is smaller compared to La3Ni2O7 due to the smaller hybridization with $d_{3z^{2}-r^{2}}$. The sign change between the bonding and antibonding $d_{3z^{2}-r^{2}}$ bands corresponds to the dominant contribution of the off-diagonal $d_{3z^{2}-r^{2}}$ component in the orbital representation, as shown in Fig. 4(c), which means that the pairing takes place mainly in the interlayer channel.
To understand the origin of the large eigenvalue around stoichiometry for Sr3Ni2O5Cl2, we plot $\lambda$ against $n[d_{3z^{2}-r^{2}}]$ in the right panel of Fig. 3(c). It can be seen from this figure that the reason for $\lambda$ being larger than La3Ni2O7 is that $n[d_{3z^{2}-r^{2}}]$ for Sr3Ni2O5Cl2 is closer to unity at stoichiometry. If we compare the two materials at the same $n[d_{3z^{2}-r^{2}}]$, $\lambda$ for Sr3Ni2O5Cl2 is suppressed compared to that of La3Ni2O7 when $\Delta n<0$, where the Fermi level intersects the nearly flat band consisting of a mixture of $d_{yz/zx}$ and $d_{3z^{2}-r^{2}}$ orbitals. This suppression becomes small when the $t_{2g}-e_{g}$ interorbital interactions are turned off, suggesting that this interorbital interaction is harmful for superconductivity especially when the $t_{2g}$ orbitals strongly hybridize with the $d_{3z^{2}-r^{2}}$ orbital near the Fermi level. This is the origin of the reduction of $\lambda$ upon hole doping.
Here, we make a brief comment on the robustness of our results against the interaction parameters. We also calculated $\lambda$ for Sr3Ni2O5Cl2 at stoichiometry using the interaction parameters evaluated with the constrained random-phase approximation (cRPA) and obtained $\lambda=0.767$. Although this value is slightly smaller than $\lambda=0.806$ shown in Fig. 4(c), it is still sufficiently higher than $\lambda$ for pressurized La3Ni2O7 at stoichiometry. The obtained interaction parameters are shown in Appendix D. We also note that, in our previous study [8], we verified that $\lambda$ for La3Ni2O7 is similar between two sets of interaction parameters: ($U,U^{\prime},J=J^{\prime}$) = (3, 2.4, 0.3) eV and those evaluated with cRPA [11].
Figure 4: (a) The superconducting gap function for Sr3Ni2O5Cl2 ($\Delta n=0.01$) at $T=0.01$ eV in the band representation, compared with (b) that for the four-orbital model of La3Ni2O7 ($\Delta n=0$). The strength of the Wannier orbital characters is shown with the radius of the color coded circles, where the weight of the $d_{3z^{2}-r^{2}}$ orbital is indicated by either green or blue, depending on whether the band energy is above or below $E=0.1$ eV. (c) Same as (a) but in the orbital representation, where the orbital-diagonal components for all $d$ orbitals and the orbital-off-diagonal component for the $d_{3z^{2}-r^{2}}$ orbital are shown.
Our theoretical analysis illustrates the effect of the mixed-anion strategy of materials design on bilayer nickelate. First of all, the tetragonal structure is stabilized in Sr3Ni2O5Cl2 as confirmed by our phonon calculation. Second, the Ni-$d_{3z^{2}-r^{2}}$ orbital energy is pushed down by the crystal field where a half of apical O2- ions is replaced with a Cl- ion. This results in two opposite effects on superconductivity as follows. The Ni-$d_{3z^{2}-r^{2}}$ orbital becomes rather closer to the half-filling in Sr3Ni2O5Cl2 than La3Ni2O7 whereas the stoichiometric Ni-$3d$ band filling are decreased in Sr3Ni2O5Cl2 ($d^{7}$) compared with La3Ni2O7 ($d^{7.5}$). The increase of the Ni-$d_{3z^{2}-r^{2}}$ band filling is beneficial for superconductivity, which was also discussed for La3Ni2O7 in our previous study [8]. On the other hand, the lowered Ni-$d_{3z^{2}-r^{2}}$ orbital energy together with the low-symmetry crystal field enhances the hybridization between the $t_{2g}$ and $e_{g}$ orbitals, which is detrimental for superconductivity. In total, we have seen that Sr3Ni2O5Cl2 is a promising new candidate of bilayer-nickelate superconductors, which can possess even higher $T_{c}$ than pressurized La3Ni2O7, at ambient pressure.
III.4 Formation energy (enthalpy)
To examine the possibility for the synthesis of Sr3Ni2O5Cl2, we calculated formation energies for some possible chemical reactions as presented in Table 2. Here, the formation energy is defined as the total energy of the right-hand side of the chemical equation relative to that for the left-hand side. Huge (negative) formation energies for reactions (a)(b) mean that the left-hand sides of (a)(b) are very unstable and unlikely the final products after the chemical reaction, while they are perhaps good starting points to promote the chemical reaction. On the other hand, we found that Sr2NiO2Cl2 + SrNiO3 and SrNiO3 + SrNiO2 + SrCl2 have similar total energies to that of the target compound Sr3Ni2O5Cl2. This result suggests that, even by starting from the relatively unstable compounds including SrO2, which was used in synthesis of similar oxychlorides, Sr2NiO2Cl2 [100], Sr2CoO3Cl, and Sr3Co2O5Cl2 [80], a synthesis can be trapped in the middle of the reaction.
However, this problem might be resolved by the high-pressure synthesis, as adopted in Ref. 100 for Sr2NiO2Cl2. Figure 5 presents that the formation enthalpy with respect to the applied pressure for chemical reactions (c)(d) that are defined in Table 2. Here, the enthalpy is defined as the sum of the total energy and the $pV$ term, where $p$ and $V$ are the pressure and the volume, respectively. We found that the formation enthalpy goes to negative even for reaction (d).
Therefore, we can expect that the material we proposed in this study, Sr3Ni2O5Cl2, can be synthesized with a help of the high-pressure synthesis. Our calculation also suggests that, under sufficiently high pressure, Sr2NiO2Cl2 can appear as impurity phases because the formation enthalpy for chemical reaction (c) is not effectively lowered by pressure as shown in Fig. 5. This result suggests that Sr2NiO2Cl2 is possibly included as an impurity phase in the high-pressure synthesis of Sr3Ni2O5Cl2 under some conditions.
Table 2: Formation energies of Sr3Ni2O5Cl2.
(a) 2SrO2 + SrCl2 + 2Ni + $\frac{1}{2}$2 $\to$ Sr3Ni2O5Cl2 --5223
(b) SrO2 + SrO + SrCl2 + 2NiO $\to$ Sr3Ni2O5Cl2 --2523
(c) Sr2NiO2Cl2 + SrNiO3 $\to$ Sr3Ni2O5Cl2 --46
(d) SrNiO3 + SrNiO2 + SrCl2 $\to$ Sr3Ni2O5Cl2 3
Figure 5: Formation enthalpy of chemical reactions (c)(d) defined in Table 2.
IV Summary
We have theoretically investigated the electronic structure of Sr3Ni2O5Cl2 and its structural stability. Our phonon calculation has shown that this compound with the $I4/mmm$ tetragonal structure is dynamically stable, i.e., does not have imaginary-frequency phonon modes, even at ambient pressure. The Ni-$d_{3z^{2}-r^{2}}$ orbital energy is pushed down by the crystal field where a half of apical O2- ions is replaced with a Cl- ion. By this effect, the Ni-$d_{3z^{2}-r^{2}}$ orbital becomes rather closer to the half-filling in Sr3Ni2O5Cl2 than La3Ni2O7 whereas the stoichiometric Ni-$3d$ band filling are decreased in Sr3Ni2O5Cl2 ($d^{7}$) compared with La3Ni2O7 ($d^{7.5}$).
By investigating superconductivity with the multi-orbital Hubbard model using the FLEX approximation combined with the linearized Eliashberg equation, we have found two features originating from the characteristic crystal field. First, interorbital interactions between the $t_{2g}$ and $e_{g}$ orbitals are detrimental for superconductivity. On the other hand, a nearly half-filled Ni-$d_{3z^{2}-r^{2}}$ orbital enhances superconductivity. In total, we have found that Sr3Ni2O5Cl2 is a promising new candidate of bilayer-nickelate superconductors, which can possess even higher $T_{c}$ than pressurized La3Ni2O7, at ambient pressure. By calculating the formation enthalpy, we have found that the high-pressure synthesis can be a good way to actually produce Sr3Ni2O5Cl2.
It is noteworthy that the bilayer band structure consisting of Ni-$d_{3z^{2}-r^{2}}$ orbitals has a quite small bandwidth with a large bilayer splitting energy in Sr3Ni2O5Cl2. Namely, the present system perhaps lies in the strong-coupling regime of the bilayer square lattice, where the interlayer superexchange interaction $J_{\perp}$ stronger than the intralayer hopping amplitudes can enhance the superconductivity [104]. This aspect can already be seen in La3Ni2O7, but is more pronounced in the present system. Thus, it is an intriguing future problem to investigate the present system from a strong-coupling picture.
Our study offers an important knowledge for materials design of bilayer-nickelate superconductors and will stimulate experimental investigation of mixed-anion bilayer-nickelate superconductors.
Acknowledgements.
We thank Yoshihiko Takano, Hiroya Sakurai, and Kazuki Yamane for fruitful discussions. This study was supported by JSPS KAKENHI Grant No. JP22K03512, JP22K04907, and JP24K01333. Part of computation was performed using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo, Japan, and Academic Center for Computing and Media Studies, Kyoto University, Japan.
Appendix A: Phonon band dispersion for La3Ni2O7
Figure 6 presents the calculated phonon dispersion for La3Ni2O7 under 0 and 20 GPa. We confirmed that the crystal structure with the $I4/mmm$ space group is unstable without applying the pressure but is stable by applying a sufficient pressure as reported in several studies [32, 42, 41].
Figure 6: Phonon band dispersion for La3Ni2O7 at (a) 0 GPa and (b) 20 GPa, respectively.
Appendix B: Phonon band dispersion for Sr3Ni2O4Cl3
Figure 7 presents the crystal structure and the calculated phonon dispersion for Sr3Ni2O4Cl3, where all the apical oxygens were replaced with chlorines. We found that the crystal structure with the $I4/mmm$ space group is unstable for this material because imaginary-frequency phonon modes appear in Fig. 7(b).
Figure 7: (a) Crystal structure and (b) phonon band dispersion for Sr3Ni2O4Cl3.
Appendix C: Electronic and phonon band structures for La0.3Sr2.7Ni2O5Cl2
Figure 8 presents the electronic band structure and the phonon dispersion for La0.3Sr2.7Ni2O5Cl2, where the La doping was handled with the virtual crystal approximation. We found that the rigid-band approximation adopted in this study was justified in Fig. 8(a) because the electronic band dispersion for La0.3Sr2.7Ni2O5Cl2 agrees well with that for Sr3Ni2O5Cl2 with a shifted Fermi energy. In addition, the absence of the imaginary-frequency phonon modes in Fig. 8(b) suggests that the stability of the $I4/mmm$ structure is kept against the La doping.
Figure 8: (a) Electronic band structure for La0.3Sr2.7Ni2O5Cl2 (blue solid lines) shown together with that for Sr3Ni2O5Cl2 with a shifted Fermi energy (black broken lines). (b) Phonon band dispersion for La0.3Sr2.7Ni2O5Cl2.
Appendix D: cRPA interaction parameters for Sr3Ni2O5Cl2
We evaluated the onsite interaction parameters for the ten-orbital model of Sr3Ni2O5Cl2 using the cRPA [105] with the projector method [106]. For this purpose, we used the plane-wave cutoff energy of 450 eV, 200 bands including occupied and unoccupied states, and an $8\times 8\times 8$ ${\bm{k}}$-mesh. Obtained interaction parameters are shown in Table 3.
Table 3: Interaction parameters of onsite Coulomb repulsion $U,U^{\prime}$ and Hund’s coupling (pair-hopping) $J$ ($J^{\prime}$) evaluated with cRPA. The orbital indices $l,m=1,2,3,4,5$ indicate the $d_{x^{2}-y^{2}},d_{3z^{2}-r^{2}},d_{xy},d_{yz},d_{zx}$ orbitals, respectively. For simplicity, we abbreviate four-orbital-index-representation of partially screened interaction integrals $V$, namely, $V_{llmm}\rightarrow U^{\prime}_{lm}$ and $V_{lmlm}(=V_{lmml})\rightarrow J_{lm}$. Note that $U^{\prime}_{lm}=U^{\prime}_{ml}$, $J_{lm}=J_{ml}$, and $J_{lm}=J^{\prime}_{lm}$. All in eV.
$l,m$ $U,U^{\prime}$ $J,J^{\prime}$
$1,1$ 3.936   —
$1,2$ 2.866 0.563
$1,3$ 3.151 0.249
$1,4$ 2.990 0.473
$1,5$ 2.990 0.473
$2,2$ 4.080   —
$2,3$ 2.683 0.460
$2,4$ 3.365 0.347
$2,5$ 3.365 0.347
$3,3$ 3.393   —
$3,4$ 2.802 0.420
$3,5$ 2.802 0.420
$4,4$ 4.088   —
$4,5$ 3.040 0.472
$5,5$ 4.088   —